# Approximating Unrelated Machine Weighted Completion Time Using Iterative Rounding and Computer Assisted Proofs

### v1

@@ Abstract
We revisit the unrelated machine scheduling problem with the weighted completion time objective. It is known that independent rounding achieves a 1.5 approximation for the problem, and many prior algorithms improve upon this ratio by leveraging strong negative correlation schemes. On each machine $i$, these schemes introduce strong negative correlation between events that some pairs of jobs are assigned to $i$, while maintaining non-positive correlation for all pairs. Our algorithm deviates from this methodology by relaxing the pairwise non-positive correlation requirement. On each machine $i$, we identify many groups of jobs. For a job $j$ and a group $B$ not containing $j$, we only enforce non-positive correlation between $j$ and the group as a whole, allowing $j$ to be positively-correlated with individual jobs in $B$. This relaxation suffices to maintain the 1.5-approximation, while enabling us to obtain a much stronger negative correlation within groups using an iterative rounding procedure: at most one job from each group is scheduled on $i$. We prove that the algorithm achieves a $(1.36+\epsilon)$-approximation, improving upon the previous best approximation ratio of $1.4$ due to Harris. While the improvement may not be substantial, the significance of our contribution lies in the relaxed non-positive correlation condition and the iterative rounding framework. Due to the simplicity of our algorithm, we are able to derive a closed form for the weighted completion time our algorithm achieves with a clean analysis. Unfortunately, we could not provide a good analytical analysis for the quantity; instead, we rely on a computer assisted proof. Nevertheless, the checking algorithm for the analysis is easy to implement, essentially involving evaluation of maximum values of single-variable quadratic functions over given intervals. Therefore, unlike previous results which use intricate analysis to optimize the final approximation ratio, we delegate this task to computer programs.
@@ 1 Introduction
$\sum_{i\in M}\Big{(}\sum_{j\in\sigma^{-1}(i)}p_{ij}w_{j}+\sum_{\{j,j^{\prime}\}\subseteq\sigma^{-1}(i):j\neq j^{\prime}}\min\{p_{ij}w_{j^{\prime}},p_{ij^{\prime}}w_{j}\}\Big{)}.$
@@
While our improvement over the prior best ratio of $1.4$ may not be substantial, we believe the significance of our contribution lies in the relaxed non-positive correlation condition and the iterative rounding framework. The algorithm performs one of two operations in each iteration, and terminates with an integral solution when no operations can be performed. In contrast, previous algorithms need to design various strong negative correlation schemes, often involving some complex procedures. In these schemes, the parameter $\xi$ in the probability $(1-\xi)x_{ij}x_{ij^{\prime}}$ that two jobs in a same group tends to be small, while our procedure guarantees $\xi=1$. Due to the simplicity of our algorithm, we are able to derive a closed form for the weighted completion time the algorithm provides with a clean analysis. Unfortunately, we could not provide a good analytical analysis for the ratio between the derived quantity and the LP cost; instead, we rely on a computer assisted proof. Nonetheless, the checking algorithm for the analysis is easy to implement, essentially involving the evaluation of maximum values of single-variable quadratic functions over given intervals. This leads to another advantage of our result: while previous results use involved analysis to optimize the approximation ratio, we delegate the tasks of optimizing the verifying the ratio to computer programs. Finally, it is worth noting two features of our algorithm. Firstly, it is the first algorithm to leverage the condition that the weights $w_{j}$’s are independent of machines. Prior algorithms do not use this condition and thus they are capable of solving the more general scheduling problem where we also have machine-dependent weights $w_{ij}$. Secondly, our algorithm is based on the configuration LP relaxation for the problem, marking the first use of this relaxation in designing improved approximation algorithms for the scheduling problem.
Overview of Our Algorithm and Analysis
For the sake of convenience, we swap job sizes and weights to obtain a scheduling instance with machine-independent job sizes $p_{j}$, but machine dependent weights $w_{ij}$. It is a simple observation that the swapping does not change the instance. We solve the configuration LP to obtain a fractional assignment $(z_{ij})_{i\in M,j\in J}$. With the relaxed non-positive correlation condition, we design our iterative rounding algorithm as follows. Let $G=(M,J,E)$ be the support bipartite graph of $z$. The volume of an edge $ij$ is defined as $p_{j}z_{ij}$, and we use $\mathrm{vol}(E^{\prime})$ to denote the total volume edges in $E^{\prime}$ for every $E^{\prime}\subseteq E$. We partition jobs into classes based on their sizes geometrically, with a random shift: a job $j$ is in $J_{k},k\in{\mathbb{Z}}$, if we have $p_{j}\in[\beta\rho^{k},\beta\rho^{k+1})$, for some $\rho>1$ and a randomly chosen $\beta\in[1,\rho)$. Our iterative rounding algorithm handles each job class $J_{k}$ separately and independently; henceforth we fix $k$. For every machine $i$, let $\delta^{k}_{i}$ be the set of edges between $i$ and $J_{k}$. We sort the edges $ij\in\delta^{k}_{i}$ in descending order of their $\frac{w_{ij}}{p_{j}}$ values, i.e., Smith ratios. We mark the first $\rho\beta^{k}$ volume the edges in this order, or all edges in $\delta^{k}_{i}$ if $\mathrm{vol}(\delta^{k}_{i})\leq\beta\rho^{k}$; the other edges in $\delta^{k}_{i}$ are unmarked. Let $\delta^{k,\mathrm{mk}}_{i}$ be the set of marked edges in $\delta^{k}_{i}$. In the actual algorithm, we may need to break some edge into two parallel edges, one marked and other unmarked. For simplicity, we assume this does not happen in the overview. We maintain a $\bar{x}_{ij}$ value for every $ij\in E$ between $M$ and $J_{k}$, initially set to $z_{ij}$. Let $\bar{G}=(M,J_{k},\bar{E})$ be the support of $\bar{x}$. A crucial property we maintain is that $\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}_{e}p_{e}$ does not change (we define $p_{e}=p_{j}$ if $e=ij$), until when $|\delta^{k,\mathrm{mk}}_{i}\cap\bar{E}|=1$ happens. Two operations respecting the property can be defined: a rotation operation over a cycle of marked edges in $\bar{E}$, and a shifting operation along a “pseudo-marked-path” in $\bar{E}$. Both operations are random and respect the marginal probabilities, and the condition that the every job is assigned to an extent of 1. As the former operation is standard, we only describe the latter. A pseudo-marked-path consists of a simple path $(j_{1},i_{1},j_{2},i_{2}\cdots,i_{t-1},j_{t})$ of edges in $\bar{E}$ between two machines $j_{1}$ and $j_{t}$, with one edge $i_{0}j_{1}$ at the beginning and one edge $j_{t}i_{t}$ in the end. The edge $i_{0}j_{1}$ may be unmarked, in which case we allow $i_{0}\in\{i_{1},i_{2},\cdots,i_{t-1}\}$, or the unique edge in $\delta^{k,\mathrm{mk}}_{i_{0}}\cap\bar{E}$. The same requirement is imposed on $j_{t}i_{t}$. Naturally, we shift $\bar{x}$-values on the path in one of the two directions randomly. Each operation will remove at least 1 edge from $\bar{G}$, and the procedure terminates when we have no cycles of marked edges or pseudo-marked-paths in $\bar{E}$. We prove that when this happens, the $\bar{x}$-vector is integral, which gives the assignment of $J_{k}$ to $M$. Now we focus on the correlation between two edges in $\delta^{k}_{i}$. Two unmarked edges have a non-positive correlation. Positive correlations may be introduced between an unmarked edge and a marked one by the shifting operation along a pseudo-marked path. It may happen that $i_{1}j_{1}$ is unmarked and $i_{1}$ is the same as some $i_{o}$ on the path. Then $i_{1}j_{1}$ and $i_{o}j_{o}$ will be positively related. However, the quantity $\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}_{e}p_{e}$ does not change; that is, $j$ is non-positively correlated with $\delta^{k,\mathrm{mk}}_{i}$ as a whole. This is where the relaxed non-positive correlation condition comes into play. As all the marked edges precede unmarked ones in the final schedule, this suffices to preserve the $1.5$-approximation ratio. The improvement over 1.5 arises from the strong negative correlation among marked edges. We maintain $\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}_{e}p_{e}\leq\beta\rho^{k}$ until $\delta^{k,\mathrm{mk}}_{i}\cap\bar{E}$ becomes a singleton. As every class-$k$ job has size at least $\beta\rho^{k}$, at most one edge in $\delta^{k,\mathrm{mk}}_{i}$ will be chosen in our final schedule, resulting in the strongest possible negative correlation within $\delta^{k,\mathrm{mk}}_{i}$. For the analysis, we derive a relatively simple closed form for the expected total weighted completion time. We can separately focus on each machine $i$, a threshold $\sigma$ on the Smith ratio, and all the jobs $j$ with $\frac{w_{ij}}{p_{i}}\geq\sigma$, denoting them as $J^{\prime}$. The term leading to the improvement over $1.5$ is precisely expressed as $\frac{1}{2}\sum_{k}\big{(}\mathrm{vol}(\delta^{k,\mathrm{mk}}_{i}\cap J^{\prime})\big{)}^{2}=\frac{1}{2}\sum_{k}\big{(}\min\big{\{}\mathrm{vol}\big{(}\delta^{k}_{i}\cap J^{\prime}),\beta\rho^{k}\big{\}}\big{)}^{2}$. While this term is simple, we still encounter difficulties in providing a good analytical analysis of our ratio. Instead, we provide a computer assisted proof with $\rho=2$. We only focus on three most important job classes $k$, disregarding the savings from others. The ratio is captured by some mathematical program with a maximization objective. Depending on whether $\mathrm{vol}(\delta^{k,\mathrm{mk}}_{i})=\mathrm{vol}(\delta^{k}_{i})$ or $\mathrm{vol}(\delta^{k,\mathrm{mk}}_{i})=\beta\rho^{k}_{i}$ for each of the three classes $k$, we further partition the program into three sub-programs. By specifying some Lagrangian multipliers for the sub-programs, we establish upper bounds on the sub-programs, and thus the original program. This results in our $(1.36+\epsilon)$-approximation ratio. A computer program verifies the validity of all the multipliers and upper bounds. The checker algorithm is easy to implement; essentially, it involves checking hundreds of quadratic function problems, each asking for the maximum value of a single-variable quadratic function over an interval.
Other Related Work
@@
Organization
The rest of the paper is organized as follows. In Section 2, we describe our iterative rounding algorithm for the scheduling problem, and show that it terminates with an integral assignment. In Section 3, we establish an upper bound on the weighted completion time given by our algorithm. In Section 4, we compare the upper bound and the LP cost to obtain our approximation ratio. It suffices to focus on a fixed machine $i$, and the set of jobs whose Smith ratios are above some threshold on $i$. We conclude with some open problems in Section 5.
@@ 2.1 Swapping Job Sizes and Weights
At the very beginning of our algorithm, we swap job sizes and weights so that in the new instance, a job $j$ has a fixed size $p_{j}$, but machine-dependent weights $w_{ij}$’s. To achieve this goal, we consider the general problem where both sizes and weights depend on the machines: We are given a weight $w_{ij}>0$ and a processing time $p_{ij}>0$ for every $i\in M$ and $j\in C$. We need to find an assignment $\phi:J\to M$ of jobs to machines. On every machine $i$, we consider the schedule of jobs $\phi^{-1}(i)$ using the Smith rule, with job processing times $(p_{ij})_{j\in\phi^{-1}(i)}$ and weights $(w_{ij})_{j\in\phi^{-1}(i)}$. This gives us a weighted completion time for machine $i$; the objective we try to minimize is the sum of this quantity over all machines $i$. A simple but useful observation we use is the following:
@@
For every machine $i$, define $\prec_{i}$ to be the total order over $J$ w.r.t the Smith ratios on machine $i$ in instance $\mathcal{I}$. That is, for every $j,j^{\prime}\in J,j\neq j^{\prime}$, $j\prec_{i}j^{\prime}$ holds if and only if $j$ should be processed before $j^{\prime}$ if they are both assigned to $i$ in the instance $\mathcal{I}$. This implies $\frac{w_{ij}}{p_{ij}}\geq\frac{w_{ij^{\prime}}}{p_{ij^{\prime}}}$; for two jobs with the same Smith ratio, we break the tie arbitrarily. Define $\prec^{\prime}_{i}$ in the same way but for the instance $\mathcal{I}^{\prime}$. Then, we can guarantee $j\prec_{i}j^{\prime}$ if and only if $j^{\prime}\prec^{\prime}_{i}j$ for every $i\in M,j,j^{\prime}\in J,j\neq j^{\prime}$. The cost of any $\phi:J\to M$ w.r.t $\mathcal{I}$ is
@@
Therefore, swapping the job sizes and weights does not change the instance. In the standard unrelated machine weighted completion time problem, the weights are machine-indepdent. By the lemma, it is equivalent to the problem where weights may depend on machines but job sizes are machine-independent. From now on, we focus on such an instance. Every job $j$ has a fixed size $p_{j}>0$, but machine-dependent weights $(w_{ij})_{i\in M}$. We remark that the swapping is only for the sake of convenience, as one could obtain an equivalent algorithm without swapping them. As in [2] and [9], we shall partition jobs into classes according to their sizes, and machine-independent sizes will ensure a global partition. Without the swapping, we have to partition the jobs according to their weights, which is incompatible with prior algorithms. Also, our algorithm crucially depends on that sizes are machine-independent after swapping. So, unlike prior algorithms, ours does not work for the general scheduling problem. We shall use $p(J^{\prime}):=\sum_{j\in J^{\prime}}p_{j}$ for every $J^{\prime}\subseteq J$ to denote the total size of jobs in $J^{\prime}$.

### v2

@@ Abstract
We revisit the unrelated machine scheduling problem with the weighted completion time objective. It is known that independent rounding achieves a 1.5 approximation for the problem, and many prior algorithms improve upon this ratio by leveraging strong negative correlation schemes. On each machine $i$, these schemes introduce strong negative correlation between events that some pairs of jobs are assigned to $i$, while maintaining non-positive correlation for all pairs.
Our algorithm deviates from this methodology by relaxing the pairwise non-positive correlation requirement. On each machine $i$, we identify many groups of jobs. For a job $j$ and a group $B$ not containing $j$, we only enforce non-positive correlation between $j$ and the group as a whole, allowing $j$ to be positively-correlated with individual jobs in $B$. This relaxation suffices to maintain the 1.5-approximation, while enabling us to obtain a much stronger negative correlation within groups using an iterative rounding procedure: at most one job from each group is scheduled on $i$.
We prove that the algorithm achieves a $(1.36+\epsilon)$-approximation, improving upon the previous best approximation ratio of $1.4$ due to Harris. While the improvement may not be substantial, the significance of our contribution lies in the relaxed non-positive correlation condition and the iterative rounding framework. Due to the simplicity of our algorithm, we are able to derive a closed form for the weighted completion time our algorithm achieves with a clean analysis. Unfortunately, we could not provide a good analytical analysis for the quantity; instead, we rely on a computer assisted proof. Nevertheless, the checking algorithm for the analysis is easy to implement, essentially involving evaluation of maximum values of single-variable quadratic functions over given intervals. Therefore, unlike previous results which use intricate analysis to optimize the final approximation ratio, we delegate this task to computer programs.
@@ 1 Introduction
$\sum_{i\in M}\Big{(}\sum_{j\in\phi^{-1}(i)}p_{ij}w_{j}+\sum_{\{j,j^{\prime}\}\subseteq\phi^{-1}(i):j\neq j^{\prime}}\min\{p_{ij}w_{j^{\prime}},p_{ij^{\prime}}w_{j}\}\Big{)}.$
@@ 1.1 Our Results
While our improvement over the prior best ratio of $1.4$ may not be substantial, we believe the significance of our contribution lies in the relaxed non-positive correlation condition and the iterative rounding framework. The algorithm performs one of two operations in each iteration, and terminates with an integral solution when no operations can be performed. In contrast, previous algorithms need to design various strong negative correlation schemes, often involving some complex procedures. In these schemes, the parameter $\xi$ in the probability $(1-\xi)x_{ij}x_{ij^{\prime}}$ that two jobs in a same group tends to be small, while our procedure guarantees $\xi=1$.
@@
Due to the simplicity of our algorithm, we are able to derive a closed form for the weighted completion time the algorithm provides with a clean analysis. Unfortunately, we could not provide a good analytical analysis for the ratio between the derived quantity and the LP cost; instead, we rely on a computer assisted proof. Nonetheless, the checking algorithm for the analysis is easy to implement, essentially involving the evaluation of maximum values of single-variable quadratic functions over given intervals. This leads to another advantage of our result: while previous results use involved analysis to optimize the approximation ratio, we delegate the tasks of optimizing the verifying the ratio to computer programs.
Finally, it is worth noting two features of our algorithm. Firstly, it is the first algorithm to leverage the condition that the weights $w_{j}$’s are independent of machines. Prior algorithms do not use this condition and thus they are capable of solving the more general scheduling problem where we also have machine-dependent weights $w_{ij}$. Secondly, our algorithm is based on the configuration LP relaxation for the problem, marking the first use of this relaxation in designing improved approximation algorithms for the scheduling problem.
1.2 Overview of Our Algorithm and Analysis
For the sake of convenience, we swap job sizes and weights to obtain a scheduling instance with machine-independent job sizes $p_{j}$, but machine dependent weights $w_{ij}$. It is a simple observation that the swapping does not change the instance. We solve the configuration LP to obtain a fractional assignment $(z_{ij})_{i\in M,j\in J}$.
With the relaxed non-positive correlation condition, we design our iterative rounding algorithm as follows. Let $G=(M,J,E)$ be the support bipartite graph of $z$. The volume of an edge $ij$ is defined as $p_{j}z_{ij}$, and we use $\mathrm{vol}(E^{\prime})$ to denote the total volume edges in $E^{\prime}$ for every $E^{\prime}\subseteq E$. We partition jobs into classes based on their sizes geometrically, with a random shift: a job $j$ is in $J_{k},k\in{\mathbb{Z}}$, if we have $p_{j}\in[\beta\rho^{k},\beta\rho^{k+1})$, for some $\rho>1$ and a randomly chosen $\beta\in[1,\rho)$.
Our iterative rounding algorithm handles each job class $J_{k}$ separately and independently; henceforth we fix $k$. For every machine $i$, let $\delta^{k}_{i}$ be the set of edges between $i$ and $J_{k}$. We sort the edges $ij\in\delta^{k}_{i}$ in descending order of their $\frac{w_{ij}}{p_{j}}$ values, i.e., Smith ratios. We mark the first $\rho\beta^{k}$ volume the edges in this order, or all edges in $\delta^{k}_{i}$ if $\mathrm{vol}(\delta^{k}_{i})\leq\beta\rho^{k}$; the other edges in $\delta^{k}_{i}$ are unmarked. Let $\delta^{k,\mathrm{mk}}_{i}$ be the set of marked edges in $\delta^{k}_{i}$. In the actual algorithm, we may need to break some edge into two parallel edges, one marked and other unmarked. For simplicity, we assume this does not happen in the overview.
We maintain a $\bar{x}_{ij}$ value for every $ij\in E$ between $M$ and $J_{k}$, initially set to $z_{ij}$. Let $\bar{G}=(M,J_{k},\bar{E})$ be the support of $\bar{x}$. A crucial property we maintain is that $\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}_{e}p_{e}$ does not change (we define $p_{e}=p_{j}$ if $e=ij$), until when $|\delta^{k,\mathrm{mk}}_{i}\cap\bar{E}|=1$ happens. Two operations respecting the property can be defined: a rotation operation over a cycle of marked edges in $\bar{E}$, and a shifting operation along a “pseudo-marked-path” in $\bar{E}$. Both operations are random and respect the marginal probabilities, and the condition that the every job is assigned to an extent of 1. As the former operation is standard, we only describe the latter. A pseudo-marked-path consists of a simple path $(j_{1},i_{1},j_{2},i_{2}\cdots,i_{t-1},j_{t})$ of edges in $\bar{E}$ between two jobs $j_{1}$ and $j_{t}$, with one edge $i_{0}j_{1}$ at the beginning and one edge $j_{t}i_{t}$ in the end. The edge $i_{0}j_{1}$ may be unmarked, in which case we allow $i_{0}\in\{i_{1},i_{2},\cdots,i_{t-1}\}$, or the unique edge in $\delta^{k,\mathrm{mk}}_{i_{0}}\cap\bar{E}$. The same requirement is imposed on $j_{t}i_{t}$. (See Definition 2.4 for the formal definition.) Naturally, we shift $\bar{x}$-values on the path in one of the two directions randomly. Each operation will remove at least 1 edge from $\bar{G}$, and the procedure terminates when we have no cycles of marked edges or pseudo-marked-paths in $\bar{E}$. We prove that when this happens, the $\bar{x}$-vector is integral, which gives the assignment of $J_{k}$ to $M$.
Now we focus on the correlation between two edges in $\delta^{k}_{i}$. Two unmarked edges have a non-positive correlation. Positive correlations may be introduced between an unmarked edge and a marked one by the shifting operation along a pseudo-marked path. It may happen that $i_{1}j_{1}$ is unmarked and $i_{1}$ is the same as some $i_{o}$ on the path. Then $i_{1}j_{1}$ and $i_{o}j_{o}$ will be positively related. However, the quantity $\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}_{e}p_{e}$ does not change; that is, $j$ is non-positively correlated with $\delta^{k,\mathrm{mk}}_{i}$ as a whole. This is where the relaxed non-positive correlation condition comes into play. As all the marked edges precede unmarked ones in the final schedule, this suffices to preserve the $1.5$-approximation ratio. The improvement over 1.5 arises from the strong negative correlation among marked edges. We maintain $\sum_{e\in\delta^{k,\mathrm{mk}}_{i}}\bar{x}_{e}p_{e}\leq\beta\rho^{k}$ until $\delta^{k,\mathrm{mk}}_{i}\cap\bar{E}$ becomes a singleton. As every class-$k$ job has size at least $\beta\rho^{k}$, at most one edge in $\delta^{k,\mathrm{mk}}_{i}$ will be chosen in our final schedule, resulting in the strongest possible negative correlation within $\delta^{k,\mathrm{mk}}_{i}$.
For the analysis, we derive a relatively simple closed form for the expected total weighted completion time. We can separately focus on each machine $i$, a threshold $\sigma$ on the Smith ratio, and all the jobs $j$ with $\frac{w_{ij}}{p_{i}}\geq\sigma$, denoting them as $J^{\prime}$. The term leading to the improvement over $1.5$ is precisely expressed as $\frac{1}{2}\sum_{k}\big{(}\mathrm{vol}(\delta^{k,\mathrm{mk}}_{i}\cap J^{\prime})\big{)}^{2}=\frac{1}{2}\sum_{k}\big{(}\min\big{\{}\mathrm{vol}\big{(}\delta^{k}_{i}\cap J^{\prime}),\beta\rho^{k}\big{\}}\big{)}^{2}$. While this term is simple, we still encounter difficulties in providing a good analytical analysis of our ratio.
Instead, we provide a computer assisted proof with $\rho=2$. We only focus on three most important job classes $k$, disregarding the savings from others. The ratio is captured by some mathematical program with a maximization objective. Depending on whether $\mathrm{vol}(\delta^{k,\mathrm{mk}}_{i})=\mathrm{vol}(\delta^{k}_{i})$ or $\mathrm{vol}(\delta^{k,\mathrm{mk}}_{i})=\beta\rho^{k}_{i}$ for each of the three classes $k$, we further partition the program into three sub-programs. By specifying some Lagrangian multipliers for the sub-programs, we establish upper bounds on the sub-programs, and thus the original program. This results in our $(1.36+\epsilon)$-approximation ratio. A computer program verifies the validity of all the multipliers and upper bounds. The checker algorithm is easy to implement; essentially, it involves checking hundreds of quadratic function problems, each asking for the maximum value of a single-variable quadratic function over an interval.
1.3 Other Related Work
@@
Organization  The rest of the paper is organized as follows. In Section 2, we describe our iterative rounding algorithm for the scheduling problem, and show that it terminates with an integral assignment. In Section 3, we establish an upper bound on the weighted completion time given by our algorithm. In Section 4, we compare the upper bound and the LP cost to obtain our approximation ratio. It suffices to focus on a fixed machine $i$, and the set of jobs whose Smith ratios are above some threshold on $i$. We conclude with some open problems in Section 5.
@@ 2.1 Swapping Job Sizes and Weights
At the very beginning of our algorithm, we swap job sizes and weights so that in the new instance, a job $j$ has a fixed size $p_{j}$, but machine-dependent weights $w_{ij}$’s. To achieve this goal, we consider the general problem where both sizes and weights depend on the machines: We are given a weight $w_{ij}>0$ and a processing time $p_{ij}>0$ for every $i\in M$ and $j\in C$. We need to find an assignment $\phi:J\to M$ of jobs to machines. On every machine $i$, we consider the schedule of jobs $\phi^{-1}(i)$ using the Smith rule, with job processing times $(p_{ij})_{j\in\phi^{-1}(i)}$ and weights $(w_{ij})_{j\in\phi^{-1}(i)}$. This gives us a weighted completion time for machine $i$; the objective we try to minimize is the sum of this quantity over all machines $i$.
A simple but useful observation we use is the following:
@@
For every machine $i$, define $\prec_{i}$ to be the total order over $J$ w.r.t the Smith ratios on machine $i$ in instance $\mathcal{I}$. That is, for every $j,j^{\prime}\in J,j\neq j^{\prime}$, $j\prec_{i}j^{\prime}$ holds if and only if $j$ should be processed before $j^{\prime}$ if they are both assigned to $i$ in the instance $\mathcal{I}$. This implies $\frac{w_{ij}}{p_{ij}}\geq\frac{w_{ij^{\prime}}}{p_{ij^{\prime}}}$; for two jobs with the same Smith ratio, we break the tie arbitrarily. Define $\prec^{\prime}_{i}$ in the same way but for the instance $\mathcal{I}^{\prime}$. Then, we can guarantee $j\prec_{i}j^{\prime}$ if and only if $j^{\prime}\prec^{\prime}_{i}j$ for every $i\in M,j,j^{\prime}\in J,j\neq j^{\prime}$.
The cost of any $\phi:J\to M$ w.r.t $\mathcal{I}$ is
@@
Therefore, swapping the job sizes and weights does not change the instance. In the standard unrelated machine weighted completion time problem, the weights are machine-indepdent. By the lemma, it is equivalent to the problem where weights may depend on machines but job sizes are machine-independent.
From now on, we focus on such an instance. Every job $j$ has a fixed size $p_{j}>0$, but machine-dependent weights $(w_{ij})_{i\in M}$. We remark that the swapping is only for the sake of convenience, as one could obtain an equivalent algorithm without swapping them. As in [2] and [9], we shall partition jobs into classes according to their sizes, and machine-independent sizes will ensure a global partition. Without the swapping, we have to partition the jobs according to their weights, which is incompatible with prior algorithms. Also, our algorithm crucially depends on that sizes are machine-independent after swapping. So, unlike prior algorithms, ours does not work for the general scheduling problem.
We shall use $p(J^{\prime}):=\sum_{j\in J^{\prime}}p_{j}$ for every $J^{\prime}\subseteq J$ to denote the total size of jobs in $J^{\prime}$.